OpenStudy (anonymous):

guys correct me if im wrong, ∫ cos 3x dx answer: =∫cos(u)1/3 dx cos(u) =∫---- dx 3 =1/3 sin (3x) final answer : = sin(3x) ----- +c 3 correct?

OpenStudy (anonymous):

wrong or not?

OpenStudy (anonymous):

@campbell_st sir

OpenStudy (campbell_st):

well I know that \[\int\limits \cos(ax) ~dx = \frac{1}{a} \sin(ax) + C\] hope that helps

OpenStudy (campbell_st):

so your answer looks correct

OpenStudy (anonymous):

thank you sir!!

ganeshie8 (ganeshie8):

One nice way to double check the answer is by simply differentiating the answer you should get back the integrand if your answer is correct

OpenStudy (anonymous):

@campbell_st is right. you've got the correct answer.